# Generalized Kummer surfaces over finite fields

### v3

@@ 4. Rigid and symplectic actions on abelian surfaces
In this section we classify finite groups with a rigid and symplectic action on an abelian surface over $k$. Let $r$ be a power of an odd prime $p$. Denote by $\operatorname{ESL}_{2}({\mathbb{F}}_{r})$ the subgroup of $\operatorname{SL}_{2}({\mathbb{F}}_{r^{2}})$ generated by $\operatorname{SL}_{2}({\mathbb{F}}_{r})$ and an element given by the diagonal matrix
@@
Suppose that $p$ divides the order of $G$, and $Q$ is a Sylow $p$–subgroup of $G$. Then one of the following cases occurs:
@@
    $\operatorname{ESL}_{2}({\mathbb{F}}_{q})$, where $q$ is a power of $p$.
@@
Remark 4.4.
In cases $(4),(5)$, and $(6)$ the algebra ${\mathbb{Q}}[G]^{\mathrm{rig}}$ is a subalgebra of $\operatorname{Mat}_{2}({\mathbb{H}}_{p})$. According to Theorem 2.2, and Corollary 3.9, there is a rigid action of the group $G$ on some supersingular abelian surface. In Section 6 we study rigid and symplectic actions in other cases.
@@
According to Lemma 2.13, any element of order prime to $p$ has a non-trivial image in $\operatorname{SL}_{2}(k)$. We have to prove the lemma for an element $g$ of order $5$ and $p=5$. The action of $g$ on the Diedonné module $M(A)$, induces an action of ${\mathbb{Z}}[\zeta_{5}]$ on $M(A)$. Note that ${\mathbb{Q}}_{5}[\zeta_{5}]$ is totally ramified over ${\mathbb{Q}}_{5}$; thus $M(A)$ is a free module of rank $1$ over the DVR $W(k)\otimes{\mathbb{Z}}_{5}[\zeta_{5}]$, and $1-g$ acts as a uniformiser. It follows that the action of $1-g$ on $M(A)/5M(A)$ is nilpotent with one-dimensional kernel. The cotangent space $T_{0}^{*}(A)$ is a two-dimensional quotient of $M(A)/5M(A)$; therefore, the action of $g$ on $T_{0}^{*}(A)$ is non-trivial. ∎
@@
The groups $D_{3}$, $D_{5}$, and $C_{5}\rtimes C_{4}$ have only one faithful representation, but the trace of an element of order $2$ is zero; therefore, the representations are not rigid. The groups $D_{3}\times Q_{8}$, $C_{4}.A_{4}$, and $C_{4}.A_{5}$ contain a subgroup $C_{2}^{2}$, and do not have rigid representations according to Remark 1.2.∎
@@
According to Lemma 3.13, the order of $G$ divides $240$, and any cyclic subgroup is small. If $p>5$, Lemma 4.5 and the Dickson Theorem 4.2 give first three cases. If $p=5$, then according to Lemma 4.5 and Dickson Theorem 4.2 we get that $G$ could be a group from cases $(1-3)$, $\operatorname{ESL}_{2}({\mathbb{F}}_{5})$, or an extension of $C_{5}$ by a cyclic group of order dividing $48$. It is straightforward to check that such extensions without non-small cyclic subgroups are either cyclic or belong to the following list: $D_{5}=C_{5}\rtimes C_{2}$, $Q_{20}$, $C_{5}\rtimes C_{4}$, and $C_{5}\rtimes C_{8}$. The groups $D_{5}$, and $C_{5}\rtimes C_{4}$ have no rigid representations according to Lemma 4.7.
Let $p=3$. Let $G_{1}$ be a normal $p$-subgroup of $G$ from Section 2.5; in particular, $G/G_{1}$ is isomorphic to a subgroup of $\operatorname{SL}_{2}(k)$. We have two cases: either $G_{1}$ is trivial, or $G_{1}\cong C_{3}$. In the first case it follows from the Dickson Theorem that $G$ is either a group from cases $(1-3)$, or is an extension of $C_{3}$ by a cyclic group of order dividing $80$. Since such an extension does not contain $C_{1}5$, its order divides $48$. As before, we get the following list of non-cyclic extensions: $D_{3}=C_{3}\rtimes C_{2}$, $Q_{12}=C_{3}\rtimes C_{4}$, and $C_{3}\rtimes C_{8}$. The group $D_{3}$ has no rigid representations according to Lemma 4.7.
In the second case, we have to consider extensions of $G_{1}\cong C_{3}$ by non-cyclic subgroups of $\operatorname{SL}_{2}(k)$ of order dividing $80$, i.e, $Q_{8}$, $Q_{16}$, $Q_{20}$, $Q_{40}$, or $Q_{80}$. Since $3$ is coprime to the orders of these groups, $H^{2}(G/G_{1},G_{1})$ is trivial; therefore, any extension is a semidirect product. The groups without non-small cyclic subgroups are $\operatorname{SL}_{2}({\mathbb{F}}_{3})\cong C_{3}\rtimes Q_{8}$, $C_{3}\times Q_{8}$, and $C_{3}\rtimes Q_{16}$.
@@
,${}_{2},C_{3},C_{5},C_{6},C_{10},C_{2}\times C_{2},C_{6}\times C_{2},C_{10}\times C_{2},S_{3}=D_{3}=\operatorname{SL}_{2}({\mathbb{F}}_{2}),D_{5},A_{4}=C_{2}^{2}\rtimes C_{3},A_{5}=\operatorname{SL}_{2}({\mathbb{F}}_{4}).$
Assume first that $G_{1}$ is cyclic. If $\bar{G}$ is abelian, $D_{3}$, or $D_{5}$, then there is a normal subgroup $H$ of odd order such that $\bar{G}/H$ is a $2$-group. According to Lemma 4.6, any extension of $G_{1}$ by $\bar{G}$ is uniquely determined by the $2$-Sylow subgroup $Q$ of this extension. Since $Q$ is either cyclic or isomorphic to $Q_{8}$ or $Q_{16}$ we get the following groups without non-small cyclic subgroups:
@@
If $G_{1}=Q_{16}$, then $\bar{G}$ is $C_{3}$ or $C_{5}$, and any such extension is trivial, because $\operatorname{Aut}(Q_{16})$ is a $2$-group; in this way we get two groups without non-small cyclic subgroups: $C_{3}\times Q_{16}$, and $C_{5}\times Q_{16}$. If $G_{1}=Q_{8}$, and $\bar{G}$ is $C_{3}$ or $C_{5}$, then either extension is trivial, or $G=Q_{8}\rtimes C_{3}\cong\operatorname{SL}_{2}({\mathbb{F}}_{3})$, because $\operatorname{Aut}(Q_{8})\cong S_{4}$.
Finally, assume that $G_{1}$ is $Q_{8}$, and the $2$-Sylow subgroup of $G$ is $Q_{16}$. If $\bar{G}$ is cyclic of even order, then any homomorphism from $\bar{G}$ to $\operatorname{Aut}(Q_{8})\cong S_{4}$ factors through $C_{2}$; therefore, either $G$ is $Q_{16}$, or contains a non-small cyclic subgroup. If $\bar{G}=C_{2}\times C_{2}$, then the order of the $2$-Sylow subgroup is greater than $16$. If $\bar{G}$ is $D_{3}$ or $D_{5}$, then any extension that contain $Q_{16}$ has a non-small cyclic subgroup.
@@ 5. Singularities of $A/G$
In this section we study resolutions of singularities of the quotients $A/G$. The next proposition is a generalization of the table from the Katsura’s paper [Ka87] for cyclic groups, and of results due to Fujiki [Fu] for $Q_{8}$, $Q_{12}$, and $\operatorname{SL}_{2}({\mathbb{F}}_{3})$.
@@
Let $A$ be an abelian surface with a rigid and symplectic action of a group $G$ over $k$. Assume that $p$ does not divide the order of $G$. Then the singularities of the quotient surface $A/G$, and the rank of the Neron–Severi group of the minimal resolution of singularities $X$ are as follows.
@@
Let $G$ be a finite group of order $n$ with a rigid action on an abelian variety $A$. Denote by $N(H)$ the number of points with the stabilizer isomorphic to $H$.
@@
    Let $G=C_{4}$, and let $A$ be a surface. Then $N(C_{4})=4$, and there are $6$ orbits of length $2$.
@@
Let $g\in G$ be an element of order $d$. Since the action is rigid, the Tate module $T_{\ell}(A)$ is a free module over the regular local ring ${\mathbb{Z}}_{\ell}[\zeta_{d}]$, where $\zeta_{d}$ acts as $g$. If $d=\ell$ is a prime, and $x$ is a fixed point of $G$, then $(\zeta_{\ell}-1)x=0$. Therefore, the norm
@@
We proved $(2)$. Parts $(3)$ and $(4)$ are easy.
We now prove $(5)$. Clearly, $A[2]({\bar{k}})$ is fixed by $C_{2}\subset C_{4}$; therefore, according to $(2)$, we have $N(C_{4})=4$, and $12$ points form $6$ orbits of length $2$. ∎
@@
Let $G=Q_{8}$, and let $H=C_{4}\subset G$ be a normal subgroup. According to Lemma 5.2.(1), we have to compute the orbits of $G$ on $A[2]({\bar{k}})$. The group $H$ has $4$ fixed points by Lemma 5.2.(5). Since the length of any non-trivial orbit is even, and the origin is fixed by $G$, either $N(G)=4$, or $N(G)=2$. Note that $Q_{8}$ contains $3$ cyclic subgroups of order $4$, and if $2$ such subgroups fix a point, then the point is fixed by $G$. In other words, either $N(C_{4})=0$, or $N(C_{4})=6$. In the first case, we have $N(C_{2})=12$, and in the second case $N(C_{2})=8$.
Let $G=Q_{16}$, and let $H=C_{8}\subset G$ be the only subgroup of order $8$. The action of $G/H$ on $A[2]({\bar{k}})/H$ is very simple, and we have $N(G)=N(C_{8})=2$, and $N(Q_{8})=2$. Moreover, since there are two conjugacy classes of $C_{4}$ in $G$, we have $N(C_{4})=4$, and $N(C_{2})=8$.
@@
  • $Q_{12}$ has $3$ cyclic $4$-Sylow subgroups, thus $N(G)=1$, $N(C_{4})=9$, $N(C_{3})=8$, and $N(C_{2})=6$;
  • $Q_{20}$ has $5$ cyclic $4$-Sylow subgroups, thus $N(G)=1$, $N(C_{5})=4$, and $N(C_{4})=15$;
  • $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ has $4$ cyclic $3$-Sylow subgroups, and a normal subgroup isomorphic to $Q_{8}$, thus $N(G)=1$, $N(Q_{8})=3$, $N(C_{3})=32$, and $N(C_{2})=12$;
  • $\operatorname{SL}_{2}({\mathbb{F}}_{5})$ has $10$ cyclic $3$-Sylow subgroups, $6$ cyclic $5$-Sylow subgroups, and $5$ Sylow subgroups isomorphic to $Q_{8}$, thus $N(G)=1$, $N(Q_{8})=15$, $N(C_{5})=24$, and $N(C_{3})=80$.
@@
There is a misprint for the cyclic group of order $8$ in [Ka87].
@@
Let $A$ be an abelian surface over a perfect field $k$ of characteristic $p>2$. Assume that there exists a rigid and symplectic action on $A$ of a cyclic group $G$ of order $5,8$, or $12$. Then $A$ is supersingular.
@@
Let $X$ be the minimal resolution of singularities of $A/G$. If $p$ does not divide the order of $G$, then, according to the Katsura Theorem 1.3, and Proposition 5.1, $X$ is a supersingular $K3$ surface. By [Ka87, Lemma 4.4], $A$ is also supersingular.
If $p$ divides the order of $G$, and $A$ is mixed or ordinary, then either $A$ is simple and $\operatorname{End}^{\circ}(A)\cong{\mathbb{Q}}(\zeta_{n})$, or $A$ is isogenous to the square of an ordinary elliptic curve $B$ and
@@
Furthermore, the slopes of $A$ are different; therefore, in the decomposition of $p{\mathbb{Z}}[\zeta_{n}]$ to the product of prime ideals there are at least two different primes. Since ${\mathbb{Z}}[\zeta_{n}]$ has only one prime ideal over $p$, when $p$ divides $n\in\{5,8,12\}$, the surface $A$ is supersingular.
@@
In the first case, according to Lemma 2.6, ${\mathbb{F}}_{p^{2}}\subset k$. A contradiction. In cases $(2)$ and $(3)$ the central simple algebra ${\mathbb{H}}$ of degree $8$ over ${\mathbb{Q}}$ contains ${\mathbb{Q}}(\zeta_{n})$ as a subfield of degree $4$ over ${\mathbb{Q}}$; therefore, the center of ${\mathbb{H}}$ is a subfield of ${\mathbb{Q}}(\zeta_{n})$, and ±$\sqrt{\pm p}\in{\mathbb{Q}}(\zeta_{n})$. This is impossible if $p$ does not divide $n$. ∎
@@
    Let $A$ be an abelian surface over ${\mathbb{F}}_{q}$ with a rigid and symplectic action of a cyclic group $G$ of order $n$ such that $(n,p)=1$. Let $x\in A({\mathbb{F}}_{q^{r}})$ be a point of degree $r$ fixed by $G$ such that $\varphi_{x}$ is surjective. Then the image of $x$ is an ${\mathbb{F}}_{q}$-point, the quotient singularity at $x$ is of type $A_{n/r-1}$, and the Frobenius action on the graph of exceptional curves is non-trivial if and only if $\zeta_{n/r}\not\in{\mathbb{F}}_{q^{r}}$.
@@
    Assume that $p>2$. Let $G=Q_{8}$, and let $A$ be an abelian surface over ${\mathbb{F}}_{q}$ with a rigid and symplectic action of $G$. Then the Frobenius action on the graph $D_{4}$ of the quotient singularity at the origin of $A$ is trivial.
@@
    Since $g^{r}\in\ker\varphi_{x}$, there exists a basis $u,v$ of $V$ such that $g^{r}v=\zeta v$, and $g^{r}u=\zeta^{-1}u$, where $\zeta\in k$ is a primitive root of unity of order $n/r$. Let $\lambda$ be a generator of $\ker\pi$. Then $\lambda(v)=\alpha u$ for some $\alpha\in k$ if and only if $\lambda(\zeta)\neq\zeta$, i.e, if and only if $\zeta\not\in{\mathbb{F}}_{q^{r}}$. In what follows we will assume that $\alpha=1$.
@@
In this section we use Corollary 3.9 to obtain a classification of finite groups that act on abelian varieties over a given finite field $k$ of characteristic $p$.
@@
The first column of the table follows from Theorem 2.2 and the fact that ${\mathbb{Q}}(\zeta_{n})\otimes_{\mathbb{Q}}{\mathbb{H}}_{p}$ represents the trivial element of the Brauer group if and only if ${\mathbb{Q}}(\zeta_{n})$ does not split in $p$.
@@
Assume that ${\mathbb{F}}_{p^{2}}$ is a subfield of a perfect field $k$, and the order of $G$ is prime to $p$ and is greater than $2$. Let $A$ be a supersingular abelian surface over $k$ with a rigid action of a finite group $G$. Then $G$ and $p$ satisfy the conditions in the column $I$. If the action is symplectic, then $G$ and $p$ satisfy the conditions in the column $II$.
If $p$ and $G$ satisfy a condition in the column $I$ of the table below, then there exists a supersingular abelian surface over ${\mathbb{F}}_{p^{2}}$ with a rigid action of a finite group $G$. If $p$ does not divide the order of $G$, then there exists a supersingular abelian surface over ${\mathbb{F}}_{p^{2}}$ with a rigid and symplectic action of a group $G$ if $p$ and $G$ satisfy the conditions in the column $II$.
$G$ $I$ $II$
@@
First, we prove several lemmas in a slightly more general situation. Let $k$ be a perfect field, and let $A$ be an abelian surface over $k$ with a rigid action of a finite group $G$. If $p$ does not divide $n$, we denote by $\bar{\zeta}_{n}\in{\bar{k}}$ a primitive root of unity of degree $n$ such that $\bar{\zeta}_{n}$ lifts to $\zeta_{n}\in W({\bar{k}})$.
@@
Assume that $k$ is finite. Let $g\in G$ be an element of order $p$. Then the image of $g$ in $\operatorname{GL}(T_{0}^{*}(A))\cong\operatorname{GL}_{2}(k)$ belongs to $\operatorname{SL}_{2}(k)$. In particular, the action of $g$ is symplectic.
@@
The index of $\operatorname{SL}_{2}(k)$ in $\operatorname{GL}_{2}(k)$ is prime to $p$. ∎
@@
Let $g\in G$ be an element of order $n\in\{3,4,5,6,8,10,12\}$ prime to $p$. Assume that $\bar{\zeta}_{n}\not\in k$. Let ${\mathbb{F}}_{q}$ be the algebraic closure of ${\mathbb{F}}_{p}$ in $k$. Then the action on $T_{0}^{*}(A)$ is symplectic if and only if $q\equiv-1\bmod n$.

### v4

@@ 4. Rigid and symplectic actions on abelian surfaces
In this section we classify finite groups with a rigid and symplectic action on an abelian surface over $k$. Let $r$ be a power of an odd prime pp. Denote by $\operatorname{ESL}_{2}({\mathbb{F}}_{r})$ the subgroup of $\operatorname{SL}_{2}({\mathbb{F}}_{r^{2}})$ generated by $\operatorname{SL}_{2}({\mathbb{F}}_{r})$ and an element given by the diagonal matrix
@@
Suppose that $p$ divides the order of $G$, and $Q$ is a Sylow pp–subgroup of $G$. Then one of the following cases occurs:
@@
    $\operatorname{ESL}_{2}({\mathbb{F}}_{q})$, where $q$ is a power of pp.
@@
Remark $4.$
In cases $(4),(5)$, and $(6)$ the algebra ${\mathbb{Q}}[G]^{\mathrm{rig}}$ is a subalgebra of $\operatorname{Mat}_{2}({\mathbb{H}}_{p})$. According to Theorems 2.2 and 1.3, there is a rigid action of the group $G$ on some supersingular abelian surface. In Section 6 we study rigid and symplectic actions in other cases.
@@
According to Lemma 2.13, any element of order prime to $p$ has a non-trivial image in $\operatorname{SL}_{2}(k)$. We have to prove the lemma for an element $g$ of order 55 and $p=5$. The action of $g$ on the Diedonné module $M(A)$, induces an action of ${\mathbb{Z}}[\zeta_{5}]$ on $M(A)$. Note that ${\mathbb{Q}}_{5}[\zeta_{5}]$ is totally ramified over ${\mathbb{Q}}_{5}$; thus $M(A)$ is a free module of rank 11 over the DVR $W(k)\otimes{\mathbb{Z}}_{5}[\zeta_{5}]$, and 1g1-g acts as a uniformiser. It follows that the action of 1g1-g on $M(A)/5M(A)$ is nilpotent with one-dimensional kernel. The cotangent space $T_{0}^{*}(A)$ is a two-dimensional quotient of $M(A)/5M(A)$; therefore, the action of $g$ on $T_{0}^{*}(A)$ is non-trivial. ∎
@@
The groups $D_{3}$, $D_{5}$, and $C_{5}\rtimes C_{4}$ have only one faithful representation, but the trace of an element of order 22 is zero; therefore, the representations are not rigid. The groups $D_{3}\times Q_{8}$, $C_{4}.A_{4}$, and $C_{4}.A_{5}$ contain a subgroup $C_{2}^{2}$, and do not have rigid representations according to Remark 1.2.∎
@@
According to Lemma 3.12, the order of $G$ divides $240$, and any cyclic subgroup is small. If $p>5$, Lemma 4.5 and the Dickson Theorem 4.2 give first three cases. If $p=5$, then according to Lemma 4.5 and Dickson Theorem 4.2 we get that $G$ could be a group from cases (13)(1-3), $\operatorname{ESL}_{2}({\mathbb{F}}_{5})$, or an extension of $C_{5}$ by a cyclic group of order dividing $48$. It is straightforward to check that such extensions without non-small cyclic subgroups are either cyclic or belong to the following list: $D_{5}=C_{5}\rtimes C_{2}$, $Q_{20}$, $C_{5}\rtimes C_{4}$, and $C_{5}\rtimes C_{8}$. The groups $D_{5}$, and $C_{5}\rtimes C_{4}$ have no rigid representations according to Lemma 4.7.
Let $p=3$. Let $G_{1}$ be a normal $p$-subgroup of $G$ from Section 2.5; in particular, $G/G_{1}$ is isomorphic to a subgroup of $\operatorname{SL}_{2}(k)$. We have two cases: either $G_{1}$ is trivial, or $G_{1}\cong C_{3}$. In the first case it follows from the Dickson Theorem that $G$ is either a group from cases (13)(1-3), or is an extension of $C_{3}$ by a cyclic group of order dividing $80$. Since such an extension does not contain $C_{1}5$, its order divides $48$. As before, we get the following list of non-cyclic extensions: $D_{3}=C_{3}\rtimes C_{2}$, $Q_{12}=C_{3}\rtimes C_{4}$, and $C_{3}\rtimes C_{8}$. The group $D_{3}$ has no rigid representations according to Lemma 4.7.
In the second case, we have to consider extensions of $G_{1}\cong C_{3}$ by non-cyclic subgroups of $\operatorname{SL}_{2}(k)$ of order dividing $80$, i.e, $Q_{8}$, $Q_{16}$, $Q_{20}$, $Q_{40}$, or $Q_{80}$. Since 33 is coprime to the orders of these groups, $H^{2}(G/G_{1},G_{1})$ is trivial; therefore, any extension is a semidirect product. The groups without non-small cyclic subgroups are $\operatorname{SL}_{2}({\mathbb{F}}_{3})\cong C_{3}\rtimes Q_{8}$, $C_{3}\times Q_{8}$, and $C_{3}\rtimes Q_{16}$.
@@
$C_{2},C_{3},C_{5},C_{6},C_{10},C_{2}\times C_{2},C_{6}\times C_{2},C_{10}\times C_{2},S_{3}=D_{3}=\operatorname{SL}_{2}({\mathbb{F}}_{2}),D_{5},A_{4}=C_{2}^{2}\rtimes C_{3},A_{5}=\operatorname{SL}_{2}({\mathbb{F}}_{4}).$
Assume first that $G_{1}$ is cyclic. If $\bar{G}$ is abelian, $D_{3}$, or $D_{5}$, then there is a normal subgroup $H$ of odd order such that $\bar{G}/H$ is a 22-group. According to Lemma 4.6, any extension of $G_{1}$ by $\bar{G}$ is uniquely determined by the 22-Sylow subgroup $Q$ of this extension. Since $Q$ is either cyclic or isomorphic to $Q_{8}$ or $Q_{16}$ we get the following groups without non-small cyclic subgroups:
@@
If $G_{1}=Q_{16}$, then $\bar{G}$ is $C_{3}$ or $C_{5}$, and any such extension is trivial, because $\operatorname{Aut}(Q_{16})$ is a 22-group; in this way we get two groups without non-small cyclic subgroups: $C_{3}\times Q_{16}$, and $C_{5}\times Q_{16}$. If $G_{1}=Q_{8}$, and $\bar{G}$ is $C_{3}$ or $C_{5}$, then either extension is trivial, or $G=Q_{8}\rtimes C_{3}\cong\operatorname{SL}_{2}({\mathbb{F}}_{3})$, because $\operatorname{Aut}(Q_{8})\cong S_{4}$.
Finally, assume that $G_{1}$ is $Q_{8}$, and the 22-Sylow subgroup of $G$ is $Q_{16}$. If $\bar{G}$ is cyclic of even order, then any homomorphism from $\bar{G}$ to $\operatorname{Aut}(Q_{8})\cong S_{4}$ factors through $C_{2}$; therefore, either $G$ is $Q_{16}$, or contains a non-small cyclic subgroup. If $\bar{G}=C_{2}\times C_{2}$, then the order of the 22-Sylow subgroup is greater than $16$. If $\bar{G}$ is $D_{3}$ or $D_{5}$, then any extension that contain $Q_{16}$ has a non-small cyclic subgroup.
@@ 5. Singularities of $A/G$
In this section we study resolutions of singularities of the quotients $A/G$. The next proposition is a generalization of the table from Katsura’s paper [Ka87] for cyclic groups and of results due to Fujiki [Fu88] for $Q_{8}$, $Q_{12}$, and $\operatorname{SL}_{2}({\mathbb{F}}_{3})$.
@@
Let AA be an abelian surface with a rigid and symplectic action of a group $G$ over $k$. Assume that $p$ does not divide the order of $G$. Then the singularities of the quotient surface $A/G$, and the rank of the Neron–Severi group of the minimal resolution of singularities $X$ are as follows.
@@
Let $G$ be a finite group of order $n$ with a rigid action on an abelian variety AA. Denote by $N(H)$ the number of points with the stabilizer isomorphic to $H$.
@@
    Let $G=C_{4}$, and let AA be a surface. Then $N(C_{4})=4$, and there are 66 orbits of length 22.
@@
Let $g\in G$ be an element of order $d$. Since the action is rigid, the Tate module $T_{\ell}(A)$ is a free module over the regular local ring ${\mathbb{Z}}_{\ell}[\zeta_{d}]$, where $\zeta_{d}$ acts as $g$. If $d=\ell$ is a prime and $x$ is a fixed point of $G$, then $(\zeta_{\ell}-1)x=0$. Therefore, the norm
@@
We proved $(2)$. The parts $(3)$ and $(4)$ are easy.
We now prove $(5)$. Clearly, $A[2]({\bar{k}})$ is fixed by $C_{2}\subset C_{4}$; therefore, according to $(2)$, we have $N(C_{4})=4$, and $12$ points form 66 orbits of length 22. ∎
@@
Let $G=Q_{8}$, and let $H=C_{4}\subset G$ be a normal subgroup. According to Lemma 5.2.(1), we have to compute the orbits of $G$ on $A[2]({\bar{k}})$. The group $H$ has 44 fixed points by Lemma 5.2.(5). Since the length of any non-trivial orbit is even, and the origin is fixed by $G$, either $N(G)=4$, or $N(G)=2$. Note that $Q_{8}$ contains 33 cyclic subgroups of order 44, and if 22 such subgroups fix a point, then the point is fixed by $G$. In other words, either $N(C_{4})=0$, or $N(C_{4})=6$. In the first case, we have $N(C_{2})=12$, and in the second case $N(C_{2})=8$.
Let $G=Q_{16}$, and let $H=C_{8}\subset G$ be the only subgroup of order 88. The action of $G/H$ on $A[2]({\bar{k}})/H$ is very simple, and we have $N(G)=N(C_{8})=2$, and $N(Q_{8})=2$. Moreover, since there are two conjugacy classes of $C_{4}$ in $G$, we have $N(C_{4})=4$, and $N(C_{2})=8$.
@@
  • $Q_{12}$ has 33 cyclic 44-Sylow subgroups, thus $N(G)=1$, $N(C_{4})=9$, $N(C_{3})=8$, and $N(C_{2})=6$;
  • $Q_{20}$ has 55 cyclic 44-Sylow subgroups, thus $N(G)=1$, $N(C_{5})=4$, and $N(C_{4})=15$;
  • $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ has 44 cyclic 33-Sylow subgroups, and a normal subgroup isomorphic to $Q_{8}$, thus $N(G)=1$, $N(Q_{8})=3$, $N(C_{3})=32$, and $N(C_{2})=12$;
  • $\operatorname{SL}_{2}({\mathbb{F}}_{5})$ has $10$ cyclic 33-Sylow subgroups, 66 cyclic 55-Sylow subgroups, and 55 Sylow subgroups isomorphic to $Q_{8}$, thus $N(G)=1$, $N(Q_{8})=15$, $N(C_{5})=24$, and $N(C_{3})=80$.
@@
There is a misprint for the cyclic group of order 88 in [Ka87].
@@
Let AA be an abelian surface over a perfect field $k$ of characteristic $p>2$. Assume that there exists a rigid and symplectic action on AA of a cyclic group $G$ of order $5,8$, or $12$. Then AA is supersingular.
@@
Let $X$ be the minimal resolution of singularities of $A/G$. If $p$ does not divide the order of $G$, then, according to the Katsura Theorem 1.4, and Proposition 5.1, $X$ is a supersingular $K3$ surface. By [Ka87, Lemma 4.4], AA is also supersingular.
If $p$ divides the order of $G$, and AA is mixed or ordinary, then either AA is simple and $\operatorname{End}^{\circ}(A)\cong{\mathbb{Q}}(\zeta_{n})$, or AA is isogenous to the square of an ordinary elliptic curve $B$ and
@@
Furthermore, the slopes of AA are different; therefore, in the decomposition of $p{\mathbb{Z}}[\zeta_{n}]$ to the product of prime ideals there are at least two different primes. Since ${\mathbb{Z}}[\zeta_{n}]$ has only one prime ideal over $p$, when $p$ divides $n\in\{5,8,12\}$, the surface AA is supersingular.
@@
In the first case, according to Lemma 2.6, ${\mathbb{F}}_{p^{2}}\subset k$. A contradiction. In cases $(2)$ and $(3)$ the central simple algebra $\mathbb{H}$} of degree 88 over $\mathbb{Q}$} contains ${\mathbb{Q}}(\zeta_{n})$ as a subfield of degree 44 over $\mathbb{Q}$}; therefore, the center of $\mathbb{H}$} is a subfield of ${\mathbb{Q}}(\zeta_{n})$, and $\sqrt{\pm p}\in{\mathbb{Q}}(\zeta_{n})$. This is impossible if $p$ does not divide $n$. ∎
@@
    Let AA be an abelian surface over ${\mathbb{F}}_{q}$ with a rigid and symplectic action of a cyclic group $G$ of order $n$ such that $(n,p)=1$. Let $x\in A({\mathbb{F}}_{q^{r}})$ be a point of degree $r$ fixed by $G$ such that $\varphi_{x}$ is surjective. Then the image of $x$ is an ${\mathbb{F}}_{q}$-point, the quotient singularity at $x$ is of type $A_{n/r-1}$, and the Frobenius action on the graph of exceptional curves is non-trivial if and only if $\zeta_{n/r}\not\in{\mathbb{F}}_{q^{r}}$.
@@
    Assume that $p>2$. Let $G=Q_{8}$, and let AA be an abelian surface over ${\mathbb{F}}_{q}$ with a rigid and symplectic action of $G$. Then the Frobenius action on the graph $D_{4}$ of the quotient singularity at the origin of AA is trivial.
@@
    Since $g^{r}\in\ker\varphi_{x}$, there exists a basis $u,v$ of $V$ such that $g^{r}v=\zeta v$, and $g^{r}u=\zeta^{-1}u$, where $\zeta\in k$ is a primitive root of unity of order $n/r$. Let $\lambda$ be a generator of $\ker\pi$. Then $\lambda(v)=\alpha u$ for some $\alpha\in k$ if and only if $\lambda(\zeta)\neq\zeta$, i.e, if and only if $\zeta\not\in{\mathbb{F}}_{q^{r}}$. In what follows, we will assume that $\alpha=1$.
@@
In this section we use Theorem 1.3 to obtain a classification of finite groups that act on abelian varieties over a given finite field $k$ of characteristic pp.
@@
The first column of the table follows from Theorem 2.2 and the fact that ${\mathbb{Q}}(\zeta_{n})\otimes_{\mathbb{Q}}{\mathbb{H}}_{p}$ represents the trivial element of the Brauer group if and only if ${\mathbb{Q}}(\zeta_{n})$ does not split at pp.
@@
Assume that ${\mathbb{F}}_{p^{2}}$ is a subfield of a perfect field $k$, and that the order of $G$ is prime to $p$ and is greater than 22. Let AA be a supersingular abelian surface over $k$ with a rigid action of a finite group $G$. Then $G$ and $p$ satisfy the conditions in the column II. If the action is symplectic, then $G$ and $p$ satisfy the conditions of the column IIII.
If $p$ and $G$ satisfy a condition in the column II of the table below, then there exists a supersingular abelian surface over ${\mathbb{F}}_{p^{2}}$ with a rigid action of a finite group $G$. If $p$ does not divide the order of $G$, then there exists a supersingular abelian surface over ${\mathbb{F}}_{p^{2}}$ with a rigid and symplectic action of a group $G$ if $p$ and $G$ satisfy the conditions of the column IIII.
$G$ II IIII
@@
First, we prove several lemmas in a slightly more general situation. Let $k$ be a perfect field, and let AA be an abelian surface over $k$ with a rigid action of a finite group $G$. If $p$ does not divide $n$, we denote by $\bar{\zeta}_{n}\in{\bar{k}}$ a primitive root of unity of degree $n$ such that $\bar{\zeta}_{n}$ lifts to $\zeta_{n}\in W({\bar{k}})$.
@@
Assume that $k$ is finite. Let $g\in G$ be an element of order pp. Then the image of $g$ in $\operatorname{GL}(T_{0}^{*}(A))\cong\operatorname{GL}_{2}(k)$ belongs to $\operatorname{SL}_{2}(k)$. In particular, the action of $g$ is symplectic.
@@
The index of $\operatorname{SL}_{2}(k)$ in $\operatorname{GL}_{2}(k)$ is prime to pp. ∎
@@
Let $g\in G$ be an element of order $n\in\{3,4,5,6,8,10,12\}$ prime to pp. Assume that $\bar{\zeta}_{n}\not\in k$. Let ${\mathbb{F}}_{q}$ be the algebraic closure of ${\mathbb{F}}_{p}$ in $k$. Then the action on $T_{0}^{*}(A)$ is symplectic if and only if $q\equiv-1\bmod n$.